OpenStudy (anonymous):

Please help URGENT!!!!!!! Graph on a number line & rewrite using absolute value notation.

OpenStudy (anonymous):

whats the problem

OpenStudy (anonymous):

Graph on a number line: b) \[\left| x \right| \ge 16\]

OpenStudy (anonymous):

c) \[\left| x \right| \le -4\]

OpenStudy (anonymous):

d) \[\left| x \right| > -7\]

OpenStudy (anonymous):

Please EXPLAIN.. thank you

OpenStudy (anonymous):

okay the answer is c @saljudieh07

OpenStudy (anonymous):

what do you mean c? I need to graph on a number line

OpenStudy (anonymous):

oh okay the way you worded is weird